### v1

# CoT-Self-Instruct: Building high-quality synthetic prompts for reasoning and non-reasoning tasks

Abstract
We propose CoT-Self-Instruct, a synthetic data generation method that instructs LLMs to first reason and plan via Chain-of-Thought (CoT) based on the given seed tasks, and then to generate a new synthetic prompt of similar quality and complexity for use in LLM training, followed by filtering for high-quality data with automatic metrics. In verifiable reasoning, our synthetic data significantly outperforms existing training datasets, such as s1k and OpenMathReasoning, across MATH500, AMC23, AIME24 and GPQA-Diamond. For non-verifiable instruction-following tasks, our method surpasses the performance of human or standard self-instruct prompts on both AlpacaEval 2.0 and Arena-Hard.
1 Introduction
The transformative rise of Large Language Models (LLMs) has initiated a substantial paradigm shift in the domain of deep learning (Zhang et al., 2023; Guo et al., 2023; Long et al., 2024). The development of such models emphasizes scale, and relies heavily on large volumes of high-quality data (Gandhi et al., 2024; Abdin et al., 2024). However, acquiring such data from human sources can often be challenging or even impractical due to factors such as high costs, data scarcity, and privacy concerns (Kurakin et al., 2023). Furthermore, several studies (Hosking et al., 2023; Singh et al., 2023; Gilardi et al., 2023) have pointed out that human-generated data, being inherently prone to biases and errors, may not always be ideal for model training or evaluation. In this context, synthetic data emerges as a viable alternative for obtaining high-quality datasets.
Synthetic data is artificially generated to replicate the characteristics and patterns of real-world data. One innovative approach to creating such data is the Self-Instruct method (Wang et al., 2022a), which utilizes LLMs themselves to generate instruction-following examples. This method begins by selecting a small set of seed instruction-following samples, which are then used to prompt LLMs to produce additional demonstrations in a similar format. Since then a number of variants have been introduced that increase the complexity of queries (Liu et al., 2023; Zeng et al., 2024), maintain semantic diversity (Ding et al., 2023), scale the synthetic data (Yuan et al., 2023), and use these methods in self-improvement loops (Yuan et al., 2024). However, a significant challenge with these approaches is to ensure the quality and effectiveness of the generated data for language model training. Overall, generating high-quality synthetic data and optimizing its use for both reasoning and non-reasoning tasks still remains insufficiently understood.
In this paper, we present Chain-of-Thought(CoT)-Self-Instruct, a method that both (i) uses reasoning to help create high quality synthetic data; and (ii) self-filters the created data to only keep the highest quality ones, see Figure 1. We show the efficacy of this approach for creating both verifiable reasoning data and non-verifiable instruction following tasks, where in both cases using Chain-of-Thought (CoT) to help generate the examples outperforms those generated without CoT. To curate high quality verifiable data we introduce Answer-Consistency, which discards examples where the CoT-Self-Instruct-generated answer does not match the majority vote solution of the LLM, with the assumption that those examples are either incorrectly labeled or too difficult. For non-verifiable data we use the recent Rejecting Instruction Preferences (RIP) (Yu et al., 2025) method which measures the quality of prompts based on the distribution of reward model scores from LLM solutions. In both cases, filtering provides further gains. For reasoning tasks, CoT-Self-Instruct generated training data outperforms Self-Instruct and existing datasets such as s1k (Muennighoff et al., 2025) and OpenMathReasoning (Moshkov et al., 2025) across MATH500, AMC23, AIME24 and GPQA-Diamond. For non-reasoning tasks, it outperforms human data from WildChat (Zhao et al., 2024) and Self-Instruct, whether filtered or not, on both AlpacaEval 2 and ArenaHard. Overall, models trained with CoT-Self-Instruct generated training data provided the best results of the methods we tested.
Figure 1: CoT-Self-Instruct. Our method first prompts LLMs to reason and generate new instructions given seed prompts, followed by automatically curating high-quality data using either Answer-Consistency for verifiable reasoning tasks, or RIP (Yu et al., 2025) for non-verifiable tasks.
2 Related Work
Synthetic Data Generation
Synthetic data is produced using algorithms (Saxton et al., 2019), generative models (Borisov et al., 2022; Meng et al., 2022), or simulations (Vezhnevets et al., 2023), rather than being directly created by humans (Liu et al., 2024). It presents a promising solution for training models, particularly in scenarios where real-world data is scarce, expensive, or difficult to obtain. Self-Instruct (Wang et al., 2022a) first proposed a framework that lets a language model be prompted with seed data as few-shot examples in order to generate new synthetic data. Such data has been used to then self-train language models, e.g. in the Self-Rewarding framework (Yuan et al., 2024). Evol Instruct (Zeng et al., 2024) proposed to increase prompt complexity by letting the language model re-write the original prompts with increasing complexity. Other specific methods of creating complex synthetic data have been proposed, for example multi-hop question answering (Lupidi et al., 2024) or difficult reasoning questions (Yuan et al., 2025) both grounded on documents. Synthetic data has also been used to help train agents (Zhao et al., 2025; Zhou et al., 2025) and tool-use models (Mekala et al., 2024), and for rewriting pre-training data (Maini et al., 2024; Nguyen et al., 2025).
Synthetic Data Selection
Data selection is a critical component for post-training with synthetic data (and for data in general). Previously, LLM training was regarded as largely dependent on the size of available training data (Mishra et al., 2021; Wei et al., 2021; Wang et al., 2022b). More recent work has revealed that training on a smaller yet higher-quality curated set of prompts tends to be more effective in improving models’ both instruction following and reasoning capabilities (Zhou et al., 2024; Chen et al., 2024; Muennighoff et al., 2025; Ye et al., 2025). In addition to preprocessing techniques such as deduplication of similar prompts using similarity metrics such as ROUGE-L similarity score (Wang et al., 2022a) or clustering (Chen et al., 2023), as language models become more powerful, prompt curation can also be facilitated by using LLMs themselves as a quality judge. Recent work studies employing powerful language models to measure the complexity, diversity and quality of instructions (Lu et al., 2023; Chen et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Li et al., 2023a). The success of RLHF for post-training (Stiennon et al., 2020; Rafailov et al., 2024) has attracted more attention to collecting large scale and high quality preference data. Most work involving preference optimization employs existing methods derived from pretraining and instruction-tuning (Touvron et al., 2023; Muennighoff et al., 2025), such as deduplication, clustering, quality classifiers or filtering heuristics. Rejecting Instruction Preferences (RIP) (Yu et al., 2025) is a recent method that gives strong performance by leveraging reward models based on the LLM responses to filter prompts. For verifiable reasoning tasks, Self-Consistency filtering (Prasad et al., 2024) has also been shown to be a high quality curation method by rejecting prompts where LLM solutions show no agreement, as the task is either incorrectly labeled or too difficult.
3 Chain-of-Thought(CoT)-Self-Instruct
(CoT)-Self-Instruct is an approach to generate high quality synthetic data for training using reasoning. We first assume access to a language model, and a small amount of high quality human-annotated seed data. We consider both verifiable reasoning domains, and non-verifiable general instruction following. Our approach involves two stages:
  1. 1.
    Synthetic Instruction Creation with Chain-of-Thought (CoT): given sample human-annotated seed instructions, we instruct the LLM to reason step by step to come up with instructions of similar complexity and domain.
  2. 2.
    Synthetic Instruction Curation: we curate the generated synthetic data to keep only high-quality instructions for self-training.
We then train LLMs using the generated high quality synthetic instructions. We describe each stage in turn.
3.1 Synthetic Instruction Creation via CoT
The process of CoT-Self-Instruct data creation starts with a small set of seed instructions as the instruction pool. Multiple instructions are sampled at random from the instruction pool, and then used to few-shot prompt a language model to generate a series of intermediate reasoning steps, followed by a new instruction. Unlike standard Self-Instruct (Wang et al., 2022a) which directly prompts the model to write new instructions given a list of seed instructions, each time we show the LLM an $N$-shot set of sample instructions, we first ask it to carefully analyze the given instructions, such as domain, complexity and purpose. After analyzing the seed instructions, reflecting on what makes them high quality prompts, the LLM is prompted to reason step by step to come up with a plan to generate a new self-contained instruction that is of similar quality and complexity as the given seed instructions, and ultimately to output the final synthetic instruction satisfying these requirements in a strict answer format.
Verifiable reasoning tasks
For reasoning tasks where there is a deterministic answer which we can compare against to generate verifiable rewards during training, we instruct the LLM to use reasoning to generate both an instruction and the verifiable target. The prompt we used for CoT-Self-Instruct on reasoning tasks is given in Figure 2.
General instruction following tasks
For tasks involving general instruction-following with open-ended responses, we direct the LLM to use reasoning to generate only the instruction, not the response itself. In these instances, later during training on this synthetic data we utilize a reward model to assess the responses, eliminating the need for a reference answer. The prompt we used for CoT-Self-Instruct on general instruction following tasks is given in Figure 3. Seed prompt pools for instruction-following typically include various different domains. When selecting few-shot prompts, for example combining prompts from storytelling and coding could result in unnatural synthetic prompts. To address this, we propose to first label seed prompts into categories, and then sample from the categories first.
Figure 2: CoT-Self-Instruct prompt generation template for verifiable reasoning tasks.
[…]
Figure 3: CoT-Self-Instruct prompt generation template for general instruction following tasks.
You are a prompt generator assistant. Your goal is to create diverse and creative synthetic prompts. Please follow the steps below to create synthetic prompts. Step 1: Carefully read #Prompt 1# and #Prompt 2#. Identify and list all the common elements between these two prompts. If no common elements are found, list the main elements from each prompt. Step 2: Develop a comprehensive plan based on the #Common Elements List# or #Main Elements List# from Step 1. This plan will guide the generation of new synthetic prompts that are similar to the original prompts. Step 3: Execute the plan step by step and provide one #Synthetic Prompt#. Please reply strictly in the following format: - Step 1 #Common Elements List# or #Main Elements List#: - Step 2 #Plan#: - Step 3 #Synthetic Prompt#: #Prompt 1#: {INSTRUCTION 1}
#Prompt 2#:
{INSTRUCTION 2}
3.2 Synthetic Instruction Curation
Even with the strongest language models, not all generated synthetic instructions are well-defined and answerable, or are effective in base model self-training. We therefore apply a curation step to select higher quality synthetic instructions from the pool of generated data for final post-training with RL.
Verifiable reasoning tasks
We propose Answer-Consistency to filter and retain only high quality data. Given the task instruction, we first instruct the LLM to generate $K$ responses and take the majority response. We then reject the data example and remove it from the training pool if the majority response does not match the target answer in the synthetic data example generated by CoT-Self-Instruct (i.e., via Figure 2). Because the target answer from CoT-Self-Instruct is generated with extensive reasoning steps during instruction writing and question-answering (Figure 2) this differs from the way a response is generated at inference time and might be more accurate. Hence, comparing if the two labels match gives an extra layer of filtering. We confirm in our experiments that Answer-Consistency is superior to standard Self-Consistency filtering (Prasad et al., 2024).
General instruction following tasks
For non-verifiable tasks, the Answer-Consistency method is not applicable. Instead, we employ the Rejecting Instruction Preferences (RIP) method as proposed by Yu et al. (2025). In this method, for a given task instruction, $K$ responses are generated, and each response is evaluated using a reward model (RM), resulting in a score for each response. The filtering process is then based on the distribution of these scores. In our setting we use the lowest score among these $K$ responses to represent the score of the synthetic prompt. We then filter the data by selecting only those prompts with the higher scores. Notably, this approach can also be applied to verifiable tasks, and we conduct experiments in that context as well.
3.3 Self-training with synthetic data
After generating the synthetic training data, we can conduct RL training on the set of generated instructions. We compare the performance of self-trained LLMs with models trained on human-annotated data and on seed instructions in reasoning and non-reasoning domains respectively. For verifiable reasoning tasks, we use GRPO (Shao et al., 2024), and for general instruction following we consider both offline DPO (Rafailov et al., 2024) and online DPO, which can perform much better, see e.g. Lanchantin et al. (2025).
4 Experimental Setup
We study the effectiveness of our synthetic prompt generation approach for reasoning and non-reasoning domains along the following two axes: synthetic prompt generation, and prompt curation.
4.1 Reasoning
Seed Instructions
We use s1k (Muennighoff et al., 2025) reasoning instructions as our seed reasoning tasks. The s1k dataset consists of 1000 high-quality, diverse and difficult reasoning prompts. To conduct self-training with verifiable rewards we select a subset of s1k consisting of 893 verifiable reasoning instructions by filtering out theorem-proving questions and only keeping those that yield a scalar, single-valued, or simple closed-form answers that can be easily verified (such as 111, $A$, $False$, $\frac{2(n-1)(n-2)}{n(n+1)}$). We then use this subset as the seed instruction pool to generate more verifiable reasoning instructions.
Prompt Generation
The CoT-Self-Instruct template is given in Figure 2, and the baseline Self-Instruct method is given in Appendix Figure 6. To evaluate how CoT-Self-Instruct compares to baselines for generating verifiable reasoning tasks, we apply these methods to Qwen3-4B-Base models, Qwen3-4B model with Think mode and Qwen3-4B model with NoThink mode (Yang et al., 2025). We use temperature = 0.7 and top-p=0.8 for Qwen3-4B-Base and Qwen3-4B (NoThink mode), and temperature = 0.6 and top-p=0.95 for Qwen3-4B (Think mode).
RLVR Training
All our reasoning experiments use GRPO training initialized from Qwen3-4B-Base with reinforcement learning from rule-based verifiable rewards (RLVR). For hyperparameters, we use a cosine learning rate scheduler with a peak value of $1e-6$ and adopt the AdamW optimizer for the policy model. We set the number of training epochs to 40 with a batch size of 128. For rollouts, we sample 16 rollouts for each prompt with temperature = 0.6 and top-p=0.95, with a maximum length of 4096 tokens. All GRPO experiments are conducted with VeRL(Sheng et al., 2024) and Math-Verify  [^1] https://github.com/huggingface/Math-Verify as verifier.
Baselines & Variations
We construct targets for Self-Instruct and CoT-Self-Instruct when building the generated instructions. As a baseline, we also train on the original s1k prompts rather than synthetic data, in that case we use the public available DeepSeek R1 (Guo et al., 2025) thinking solution from simplescaling/s1K-1.1 to build targets. We also compare to training on OpenMathReasoning which consists of 10k prompts (Moshkov et al., 2025) with publicly available solutions by DeepSeek-R1 and QwQ-32B. We also explore some alternative ways of filtering data or constructing targets, explored as variations on our main experiments:
  • Self-Consistency filtering: generating responses $K$ times with random seeds and then selecting the majority-voted answer as the target or rejecting the example if the majority answer receives fewer votes than a given threshold (50% in our main experiments).
  • RIP filtering: we use the infly/INF-ORM-Llama3.1-70B (Minghao Yang, 2024) RM.
  • Best-of-K targets: constructing targets by selecting the highest scored answer out of $K$ responses using INF-ORM-Llama3.1-70B RM.
Evaluation
We evaluate using Math500 (Hendrycks et al., 2021; Lightman et al., 2023), AIME 2024, AMC 23, and GPQA Diamond (Rein et al., 2024). We use temperature 0.6 and top-p 0.95 to generate predictions. For each problem we generate N = 16 solutions and report average accuracy.
4.2 Non-verifiable instruction following
Seed Instructions
We use the Wildchat-RIP-Filtered-by-8b-Llama dataset [^2] https://huggingface.co/datasets/facebook/Wildchat-RIP-Filtered-by-8b-Llama which includes 4k high-quality prompts filtered from 20k raw wildchat prompts as our seed prompts. Unlike in some reasoning tasks, this data includes a variety of different domains. Combining, for instance, prompts from storytelling and coding as few-shot examples could result in generating unnatural synthetic prompts. To address this, we categorized all seed data into 8 distinct categories. During sampling, we select 2 seed prompts from the same category to serve as few-shot prompts. Our seed data spans 8 categories: Writing & Storytelling, Technical & Programming, Creative & Design, Data & Analysis, Education & Research, Communication & Support, Business & Marketing, and Miscellaneous.
Prompt Generation
We evaluate how CoT-Self-Instruct compares to baselines for generating non-verifiable instruction-following tasks by applying these methods using LLama 3.1-8B-Instruct. The CoT-Self-Instruct template is given in Figure 3, and the baseline Self-Instruct method is given in Figure 4. We also experiment with a prompt that lies between the two methods by generating a short rather than long CoT, given in Figure 5. For RIP filtering, we use the Athene-RM-8B model [^3] https://huggingface.co/Nexusflow/Athene-RM-8B. reward model over 32 responses.
DPO Training
We train via DPO starting from LLama 3.1-8B-Instruct, leveraging the fairseq2 library (Balioglu, 2023). We use a batch size of 646464 and learning rate of $1e{-6}$ with dropout rate of 0.0 and a $\beta$ value of 0.1 throughout the experiments. For each prompt, we generate 64 responses. These responses are then annotated with Athene-RM-8B to select pairs. Compared to human prompts, our synthetic prompts tend to be more complex, resulting in longer average response lengths, which can lead to length explosion. During DPO training, the evaluation judge often favors longer responses, potentially causing response lengths to increase over time (Yuan et al., 2024). To mitigate this issue, we adopted the approach outlined by Wu et al. (2024), which involves combining the reward score with length information to determine the preferred response. This method ensures that shorter responses are selected when scores are similar. We applied a length normalization coefficient of 0.2 for the length-normalized reward. This is applied for all methods, in each case sampling 5k DPO pairs.
Online DPO training
We also assess online DPO by following the training settings described by Lanchantin et al. (2025). We use the default sampling parameters (temperature=1.01.01.0, top-p=1.01.01.0) to generate exploration rollouts. We train models using the fairseq2 library (Balioglu, 2023), where model inference is performed with the vllm library (Kwon et al., 2023).
Evaluation
To evaluate the helpfulness and quality of responses, we employ AlpacaEval 2.0 (Li et al., 2023b; Dubois et al., 2024) and Arena-Hard (Li et al., 2024b; a). These are robust instruction-following benchmarks that show a strong correlation with user preferences. Originally, AlpacaEval used GPT-4 Preview (11/06) as the judge, while Arena-Hard utilized GPT-4.1 for its leaderboard. However, since we do not have access to these specific OpenAI API versions, we conduct our tests using two alternative (and newer) judges: GPT-4-turbo and GPT-4o. For generating predictions, we set the decoding temperature to 0.6 and the top-p to 0.9, aligning with the commonly used values of the seed model in our study. Our validation set used for checkpoint selection is based on a held-out set of 470 examples, comprising 253 validation examples from Li et al. (2023a) and 218 Evol-Test set examples from Xu et al. (2023).
5 Experimental Results
Our main results are given in Table 1 for reasoning tasks and Table 2 for non-reasoning tasks. Various other variations and ablations are given in the Appendix.
5.1 Reasoning Tasks
Synthetic instructions generated by CoT-Self-Instruct outperform Self-Instruct
In Table 1 where Qwen3-4B-Base models are GRPO trained on Qwen3-4B generated prompts and responses, CoT-Self-Instruct achieves an average accuracy of 53.0%, outperforming Self-Instruct which yields 49.5% (both without data filtering). As shown in Appendix Table 8 similar trends are observed when training on Qwen3-4B-base, rather than Qwen3-4B, generated targets.
Filtered CoT-Self-Instruct outperforms filtered Self-Instruct
Applying filtering methods to both CoT-Self-Instruct and Self-Instruct improves both methods, despite the overall amount of training data decreasing, see Table 1. That is, it is better to have high quality synthetic data than more data that is lower quality. However, we find that CoT-Self-Instruct still maintains its advantage over Self-Instruct, whichever filtering method is used. E.g. with Self-Consistency Filtering Self-Instruct improves from 49.5% $\rightarrow$ 53.6%, while CoT-Self-Instruct improves from 53.0% $\rightarrow$ 55.1%. Similar findings are observed with RIP filtering as well. We find empirically that the optimal filtering criteria is to filter out prompts with lower than 50% RIP-Scores.
Table 1: CoT-Self-Instruct results on reasoning tasks, comparing to baselines, fine-tuning Qwen3-4B-Base with GRPO. For Self-Instruct and CoT-Self-Instruct the synthetic data (including targets) is constructed with Qwen3-4B. We report pass@1 averaged over 16 seeds. CoT-Self-Instruct generates synthetic data that outperforms existing prompt training sets and the Self-Instruct method, particularly when applying our data filtering methods.
MATH AIME AMC GPQA
# Train 500 24 23 Diamond Avg. $\uparrow$
Qwen3-4B-Base (Zero-Shot) - 67.4 10.6 42.0 24.2 36.1
s1k Prompts + (R1) Gold Label 893 68.6 18.5 51.3 40.1 44.6
OpenMathReasoning Prompts + Gold Label 10,000 79.0 13.3 62.5 35.4 47.5
Self-Instruct 5000 81.1 16.3 58.1 42.5 49.5
+ Self-Consistency Filter 3467 83.6 18.5 68.5 44.1 53.6
+ RIP Filter 2254 84.5 21.2 65.9 45.5 54.5
CoT-Self-Instruct 5000 84.9 20.4 62.2 44.4 53.0
+ Self-Consistency Filter 4034 85.2 22.5 67.8 44.9 55.1
+ RIP Filter 2419 85.7 24.4 70.5 44.4 56.2
+ Answer-Consistency Filter 2926 86.5 24.6 72.3 45.5 57.2
+ Answer-Consistency Filter (more data) 10,000 86.7 26.7 73.8 47.4 58.7
Table 2: CoT-Self-Instruct results on general instruction following tasks, comparing to baselines, fine-tuning LLama 3.1-8B-Instruct with offline and online DPO. CoT-Self-Instruct generates synthetic data that outperforms human-written prompt training sets and the Self-Instruct method, particularly when applying our data filtering methods. Both AlpacaEval 2 and ArenaHard are evaluated with two kinds of judge: GPT-4 Turbo and GPT-4o, with similar conclusions.
Training AlpacaEval LC Winrate ArenaHard Score
Method GPT-4 Turbo GPT-4o GPT-4 Turbo GPT-4o Avg. $\uparrow$
LLama 3.1-8B-Instruct DPO 27.3 21.3 32.0 27.8 27.1
Human prompts (WildChat) DPO 49.1 43.0 52.7 42.6 46.8
+ RIP Filter DPO 57.6 44.5 59.1 41.7 50.7
Self-Instruct DPO 52.9 46.0 51.8 39.2 47.4
+ RIP Filter DPO 55.2 46.1 55.6 39.5 49.1
CoT-Self-Instruct DPO 58.5 48.6 62.0 46.7 53.9
+ RIP Filter DPO 63.2 49.4 60.2 45.8 54.7
Human prompts (Wildchat) Online DPO 80.1 62.7 64.4 45.5 63.1
CoT-Self-Instruct + RIP Online DPO 83.2 68.7 67.3 49.3 67.1
High quality synthetic prompts generated by CoT-Self-Instruct significantly outperform seed instructions and other publicly available reasoning prompts
CoT-Self-Instructions outperform s1k, see Table 1, where models trained on 2926 filtered CoT-Self-Instructions achieve 57.2%. This is much higher than the 44.6% achieved with s1k prompts using R1 labels (and s1k results are even lower, 43.8%, with Qwen3-4B labels, see Appendix Table 5). Filtering CoT-Self-Instruction to the same training size as s1k yields 54.2%, still significantly higher, see Appendix Table 3. These results also outperform using 10K OpenMath-Reasoning instructions with gold labels, which gives 47.5%. Increasing the CoT-Self-Instruction with Answer-Consistency filtering data to 10k improves results further with an average of 58.7%. Overall, CoT-Self-Instruct with Answer-Consistency filter gives the best performance of all existing datasets or synthetic data construction methods tested.
5.2 Non-verifiable instruction following tasks
Synthetic instructions generated by CoT-Self-Instruct outperform Self-Instruct
For non-reasoning tasks, allowing the model to create a plan beforehand with CoT-Self-Instruct also significantly enhances the quality of synthetic data, see Table 2. Averaged over AlpacaEval 2 and ArenaHard, CoT-Self-Instruct achieves an average of 53.9 vs. Self-Instruct’s 47.4, both without filtering and training with DPO. We also observe that asking for longer CoT reasoning chains provides more gains that shorter CoTs, see Appendix Table 11, further emphasizing the need for reasoning when producing synthetic data.
RIP Filtering improves CoT-Self-Instruct results further
Applying the RIP filter to each method we find it to be effective across all types of synthetic generation methods tested. This boosts the CoT-Self-Instruct results from 53.9 $\rightarrow$ 54.7. RIP also improves Self-Instruct as well, from 47.4 $\rightarrow$ 49.1, but still underperforming CoT-Self-Instruct. We can also apply RIP filtering to human prompts from WildChat in a similar manner. In this case we actually see a larger boost, from 46.8 $\rightarrow$ 50.7. We attribute this to human data being relatively noisy compared to synthetic data, which can make filtering more important.
High quality synthetic prompts generated by CoT-Instruct significantly outperform human prompts
Our best performing DPO-trained model is achieved by using CoT-Self-Instruct with RIP data filtering, yielding 54.7. This outperforms LLama 3.1-8B-Instruct (27.1) or training on human prompts from WildChat with (46.8) or without RIP data filtering (50.7). We also performed experiments with online DPO, which improved results further. In that setting human prompts from WildChat obtain 63.1 while CoT-Self-Instruct+RIP obtains 67.1. Overall, we find CoT-Self-Instruct with RIP filtering to yield the best performance over all existing datasets or synthetic data construction methods tested.
6 Conclusion
In this paper, we propose CoT-Self-Instruct, a synthetic data creation and curation pipeline that instructs LLMs to plan and reason to come up with new synthetic prompts given seed instructions, and then filters them for quality, either using Answer-Consistency for verifiable tasks or RIP filtering when they are not verifiable. We show that applying our method improves models’ abilities in both the reasoning and non-reasoning domains by creating high quality synthetic instructions for RL training, surpassing existing seed human-annotated instructions and public training sets on challenging benchmarks.